Shift summary for 10/12/99 Day Shift

Ashkan_Alavi-Harati (alavi@fnal.gov)
Tue, 12 Oct 1999 15:50:24 -0500

Shift Summary for Day shift, 12/Oct/99

On Shift: Emmanuel, Ashkan

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 6.5
Hours of E799 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: drt_fresh_start, changing tapes. Accelerator
problems.

========================================================================
Runs ended this shift:

Run 14823 - BEAM_799 - 16.6M events
Run 14824 - BEAM_799 - 4.1M events

Tapes brought to FCC this shift:
(including tapes for online split)

UP*67-69,UPS23

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Relatively smooth E799 data taking. Some DAQ problems and instable spill
structure.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14823 176 07:00 Take over the ongoing run. Beam just got
back after a couple of empty spills.

07:40 on_split tape UPF016 mounted on ktev1.

240 08:30 Ped problem, channel 997 not illuminated.

252 08:50 Stop the trigger, initialize pipeline.
The laser illumination problem is fixed.

09:00 We noticed that part of DAQ monitoring is
not updating, can not see the percentage
of the filled tapes.

10:21 on_split tape UPM014 mounted on ktev2.

11:50 MCR calls that they will try to fix the
spill structure.

394 11:53 Took advantage of the low intensity to
stop the run and do a drt_fresh_start.
12:00 Can not stop the run. Tried to save the
histograms unsuccessfully, damp server is
dead. drt_fresh_start. Also we changed
tapes. Installed new XT's in the DB.

14824 1 12:20 Start the run. The DAQ monitoring screen
is fixed. Back to the normal intensity.

12:25 on_split tape UPLA03 mounted on ktev3.

18 12:45 A couple of horrible spills, TRD planes
4 and 5 tripped. Call MCR, they said the
experts who have been working on the
problem are off to lunch! Will reset
things for now.

23 12:52 TRD's are back on.

69 13:52 No beam for the last 5 spills or so. TeV
B0 low beta frig permit toggle, est. 20
min.

74 14:01 Beam is back.

100 14:35 Quench at F3, no estimate.

102 14:39 MCR wants to control NM2SR locally, in
order to examine the power supply on that
magnet. Will call back when done.

125 15:10 Stop the run.

15:21 on_split tape UPT005 mounted on ktev4.

15:30 MCR calls, finished with the NM2SR power
supply.
========================================================================
Follow up Notes on Previous Shift

TRD planes 4 and 5 tripped off once as a result of a very high intensity
spill, otherwise stable during this shift.

SDAQ: NM2SR readback is a real problem. Rick Ford looked at the history of
the readings and the values has indeed been fluctuating recently. He
called the MCR. They will check the power supply

========================================================================
Items to Follow up on Next Shift

Keep an eye on the readback of NM2SR.

Known problems
--------------
Online histos: Banana 1X-3 (sometimes 1X-2), RC10 timing,
DC planes 9&10 inefficiency in beam region,
DC4Y intrinsic efficiency low, DC3X high SOD rms,
CsI 961 comparator errors, CsI 2310 7% gain drift.